OpenStudy (anonymous):

a bullet of mass 5g is fired using a gun of 2 kg. the velocity of the bullet is 36km/hr. what is the velocity of the gun?

OpenStudy (anonymous):

Conservation of momentum: Initially both gun and bullet are at rest thus momentum P=0 After words momentum = mass bullet * velocity + mass gun*velocity of gun = 0 since momentum is conserved. Thus mass bullet * velocity of bullet = - mass gun * velocity of gun velocity of gun = - mass bullet * velocity of bullet / mass gun v_gun= - 5g*36kg/hr /2000g= -0.09 km/hour negative sign indicates in opposite direction of bullet

OpenStudy (anonymous):

By newton's third law and law of conservation of momentum we can have both the forces balancing hence \[5*10^{-3}*(36*5/18)=(2*velocity of gun)\] velocity of gun=0.025m/s

OpenStudy (anonymous):

5 * 35 = 2000 * v

OpenStudy (anonymous):

-0.09Km/h